OpenStudy (anonymous):

With regards to Volume of Revolution for f(x) = sec^2(x), I need to integrate sec^4(x). Which is best approach - substitution or by parts?

OpenStudy (anonymous):

once you have sec^n(x), where n > 3, you'll have to use parts

OpenStudy (anonymous):

but we already had reduction formula, so it isn't that bad

OpenStudy (anonymous):

Thanks but the reduction formula isn't part of this module's content so I have to use ither parts or substitution

OpenStudy (anonymous):

Is the trick to treat it as (sec^2(x))(tan^2(x) +1)?

OpenStudy (anonymous):

substitution won't work, because one term isn't the derivative of the others. So you'll have to break it up as sec^2(x) sec^2(x), use part. u = sec^2(x), dv = sec^2(x)

OpenStudy (anonymous):

Ok so I get \[\sec ^{2}x tanx - \frac{ 2\tan ^{3}x }{3}\] This seems a little complicated compared to the rest of the examples. Going back to the Volume of Revolution, am I correct in saying that to do \[\pi \int\limits_{b}^{a} (\sec^{2}x)^{2}\] I need to to find \[\pi \int\limits_{b}^{a} \sec^{4}x\]?

OpenStudy (anonymous):

parts

OpenStudy (anonymous):

I used parts to get the top equation

OpenStudy (anonymous):

so you did u = sec^2x dv = sec^2(x) du = 2sec^2x tanx u = tan^2(x) ?

OpenStudy (anonymous):

u = tan(x) , typo

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Must've gone wrong or need further simplifying

OpenStudy (anonymous):

sec^2(x) tan(x) - 2∫tan(x) sec^2(x) dx is what you should have had

OpenStudy (anonymous):

I had sec^2(x) tan(x) - 2∫tan^2(x) sec^2(x) dx

OpenStudy (anonymous):

sec(x) sec(x) tan(x) u = sec(x) du = sec(x) tan(x) ∫u du = (1/2) u^2 = (1/2) sec^2(u) so 2 [(1/2) sec^2(x)] = sec^2(x)

OpenStudy (anonymous):

it's tan(x) sec^2(x), not tan^2(x) sec^2(x)

OpenStudy (anonymous):

sec^2(x) tan(x) - sec^2(x) is the final answer

OpenStudy (anonymous):

or sec^2(x) (tan(x) - 1)

OpenStudy (anonymous):

ahhhh..... I think I did something wrong O.O

OpenStudy (anonymous):

ok, let me back up u = sec^2(x), dv = sec^2(x) du = 2 sec(x) sec(x) tan(x), v = tan(x) so, sec^2(x) tan(x) - 2 ∫ sec(x) sec(x) tan(x) dx looks good?

OpenStudy (anonymous):

u = sec(x) du = sec(x) tan(x) then ∫u du = (1/2) u^2 what the... I did right :/

OpenStudy (anonymous):

Yeh ok, I can see that. But is there another way of writing sec^2x in terms of sin, cos or tan so I can input the limits?

OpenStudy (anonymous):

ahahah, I did it it wrong :/ should have been -2∫sec^2(x) tan^2(x) dx XD

OpenStudy (anonymous):

so you were right :D

OpenStudy (anonymous):

I was just looking back through it and getting confused! :D

OpenStudy (anonymous):

so, 2∫sec^2(x) (sec^2(x) - 1) dx 2∫sec^4 dx - 2sec^2(x) dx add 2 sec^4(x) for both sides, and integrate the last term 3∫sec^4(x) = sec^2(x) tan(x) - 2∫sec^2(x) dx 3 ∫sec^4(x) = sec^2(x) tan(x) - 2tan(x) = (1/3) sec^2(x) tan(x) - (2/3) tan(x) <---

OpenStudy (anonymous):

apparently still wrong. Must have had algebraic mistake somewhere :/

OpenStudy (anonymous):

I wonder if I'm accidentally making it too complicated. The full question is: Find the volume of the solid of revolution obtained when the graph of f(x) = sec^2x from -pi/4 to pi/3 is rotated about the x axis

OpenStudy (anonymous):

ah found it. Off by a minus sign. Should have been (1/3) sec^2(x) tan(x) + (2/3) tan(x) ^ plus, not minus

OpenStudy (anonymous):

I know the answer from Mathcad and Wolfram Alpha to be 15.072. Before multiplying by pi it should be 4.797

OpenStudy (anonymous):

just evaluate (1/3) sec^2(x) tan(x) + (2/3) tan(x) from -pi/4 to pi/3

OpenStudy (anonymous):

Ok thanks - nearly there! How do I evaluate a value with sec^2x (or indeed secx) Is it just that it's 1/cosx?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Brilliant thanks - I'll give that a go

OpenStudy (anonymous):

and multiply by pi will give the the answer

OpenStudy (anonymous):

Got it, fantastic! Thanks very much for your help!

OpenStudy (anonymous):

yw